# Vision-based Manipulation from Single Human Video with Open-World Object Graphs

### v1

@@ 6 Conclusions
In this paper, we investigate the problem of learning robot manipulation from a single human video in the open-world setting, where a robot must learn to manipulate novel objects from one video demonstration. To tackle this problem, we introduce ORION, an algorithm built on object-centric priors. Our results show that given a single human video, ORION is able to construct a policy that generalizes over the following four dimensions: visual backgrounds, camera angles, spatial layouts, and the presence of new object instances.
Limitations: We consider the task goals to be described by contact states so that we naturally avoid the ambiguities introduced when considering spatial relations, such as placing items next to an object. How to infer human intentions while clearing the inherent ambiguities in videos is a future direction to explore.
We have also assumed two constraints on how videos are captured: the camera needs to be stationary and include RGB-D data. In reality, most videos in everydayscenarios are taken while cameras are moving ande in RGB. Thus, a promising future direction is to investigate how to build a model that can reconstruct the dynamic scenes from a moving RGB camera, where the desired model can estimate the geometry of both static and moving objects in the scenes while making sure the scale of reconstructed scenes and objects matches the real world.
Furthermore, ORION establishes the correspondence between objects from demonstration and rollout using global registration of the point clouds. Such correspondence relies solely on the geometry of objects, which may suffer from ambiguities when the object shapes are symmetric and the correspondence relies on the texture information. A future direction is to incorporate both semantic and geometric information of the objects to establish object correspondence.
@@ Acknowledgments
We would like to thank Rutav Shah, Jiayuan Mao, and Fangchen Liu for the helpful discussions. This work has taken place in the Robot Perception and Learning Group (RPL) and Learning Agents Research Group (LARG) at UT Austin. RPL research has been partially supported by the National Science Foundation (FRR2145283, EFRI-2318065) and the Office of Naval Research (N00014-22-1-2204). LARG research is supported in part by NSF (FAIN-2019844, NRT-2125858), ONR (N00014-18-2243), ARO (E2061621), Bosch, Lockheed Martin, and UT Austin’s Good Systems grand challenge. Peter Stone serves as the Executive Director of Sony AI America and receives financial compensation for this work. The terms of this arrangement have been reviewed and approved by the University of Texas at Austin in accordance with its policy on objectivity in research.
@@ Appendix A Additional Technical Details
$\mathcal{G}.vo_{i}$ Object Node 3D point cloud of an object.
$\mathcal{G}.vh$ Hand Node Hand mesh and locations of the thumb and index finger.
$\mathcal{G}.vp_{ij}$ Point Node A trajectory of a TAP keypoint between two keyframes, recorded in xyz positions.
@@
Table 1: Data Structure of an OOG. For a given OOG $\mathcal{G}=(\mathcal{V},\mathcal{E})$, it has $\mathcal{V}=\{\mathcal{G}.vo_{i}\}\cup\{\mathcal{G}.vh\}\cup\{\mathcal{G}.vp_{ij}\}$, and =$\mathcal{E}=\{\mathcal{G}.eo_{ik}\}\cup\{\mathcal{G}.eh_{i}\}\cup\{\mathcal{G}.ep_{ij}\}$.
@@ A.2 Implementation Details
Object localization at test time. When we localize objects at test time, there could be some false positive segmentation of distracting objects. Such vision failures will prevent the robot policy from successfully executing actions. To exclude such false positive object segmentaiton, we use Segmentation Correspondence Model (SCM) from GROOT [11], where SCM filters out the false positive segmentation of the objects by computing the affinity scores between masks using DINOv2 features.
@@ Appendix B System Setup
Details of camera observations. As mentioned in Section 4, we use an iPad with a TrueDepth camera for collecting human video demonstrations. We use an iOS app, Record3D, that allows us to access the depth images from the TrueDepth camera. We record RGB and depth image frames in sizes $1920\times 1080$ and $640\times 480$, respectively. To align the RGB images with the depth data, we resize the RGB frames to the size $640\times 480$. The app also automatically records the camera intrinsics of the iPhone camera so that the back-projection of point clouds is made possible.

### v2

@@ 6 Conclusions
In this paper, we investigate the problem of learning robot manipulation from a single human video in the open-world setting, where a robot must learn to manipulate novel objects using one video demonstration. To tackle this problem, we introduce ORION, an algorithm built on object-centric priors. Our results show that given a single human video, with or without depth, ORION is able to construct a policy that generalizes over the following four dimensions: visual backgrounds, camera angles, spatial layouts, and the presence of new object instances.
Limitations: We consider task goals to be described by contact states, thereby naturally avoiding the ambiguities introduced when considering spatial relations, such as placing items next to an object. Inferring human intentions while clearing the inherent ambiguities in videos is a future direction to explore. We have also assumed that demonstration videos must be captured with a stationary camera. In reality, most videos in everyday scenarios contain a moving camera. Thus, a promising future direction is to investigate how to build a model that can reconstruct dynamic scenes from a moving camera, where the desired model can estimate the geometry of both static and moving objects in the scenes while ensuring the scale of the reconstructed scenes and objects matches the real world. As vision models continue to improve, stronger perceptual back-ends can be integrated into our framework with minimal modification, further boosting success rates and enabling a broader range of tasks.
@@ Acknowledgments
We would like to thank Rutav Shah, Jiayuan Mao, Fangchen Liu, Bowen Wen, and Stan Birchfield for helpful discussions. This work has taken place in the Robot Perception and Learning Group (RPL) and Learning Agents Research Group (LARG) at UT Austin. RPL research has been partially supported by the National Science Foundation (FRR2145283, EFRI-2318065) and the Office of Naval Research (N00014-22-1-2204). LARG research is supported in part by the National Science Foundation (FAIN-2019844, NRT-2125858), the Office of Naval Research (N00014-24-1-2550), Army Research Office (W911NF-17-2-0181, W911NF-23-2-0004, W911NF-25-1-0065), DARPA (Cooperative Agreement HR00112520004 on Ad Hoc Teamwork), Lockheed Martin, and Good Systems, a research grand challenge at the University of Texas at Austin. The views and conclusions contained in this document are those of the authors alone. Peter Stone serves as the Chief Scientist of Sony AI and receives financial compensation for that role. The terms of this arrangement have been reviewed and approved by the University of Texas at Austin in accordance with its policy on objectivity in research.
@@ Appendix A Additional Technical Details
$\mathcal{G}.vo_{i}$ Object Node 3D point cloud of an object (+ object mesh if RGB-only).
$\mathcal{G}.vh$ Grasp Node Grasp locations on object, calculated through hand mesh finger locations / object mesh heuristics.
$\mathcal{G}.vp_{ij}$ Point Node A trajectory of a TAP keypoint (RGB-D) or 6-DOF object pose (RGB-only) between two keyframes.
@@
Table 1: Data Structure of an OOG. For a given OOG $\mathcal{G}=(\mathcal{V},\mathcal{E})$, it has $\mathcal{V}=\{\mathcal{G}.vo_{i}\}\cup\{\mathcal{G}.vh\}\cup\{\mathcal{G}.vp_{ij}\}$, and $\mathcal{E}=\{\mathcal{G}.eo_{ik}\}\cup\{\mathcal{G}.eh_{i}\}\cup\{\mathcal{G}.ep_{ij}\}$.
@@ A.2 Implementation Details
Object localization at test time. When we localize objects at test time, there could be some false positive segmentation of distracting objects. Such vision failures will prevent the robot policy from successfully executing actions. To exclude such false positive object segmentationt, we use Segmentation Correspondence Model (SCM) from GROOT [11], where SCM filters out the false positive segmentation of the objects by computing the affinity scores between masks using DINOv2 features.
@@
FoundationPose implementation. The off-the-shelf FoundationPose implementation does not support RGB-only inputs. To adapt it for our use case, we modify the implementation to take in a tuple $(l,h)$ where $l$ and $h$ represent the lower and upper bounds of the distance from the camera to the object, respectively. We estimate this upper and lower bound by applying monocular depth estimation on the object, calculating the median depth over all pixels belonging to the object, and setting l=med0.15l=med-0.15 and $h=med+0.15$. We then create pose hypotheses centered at regular intervals between $l$ and $h$. These pose hypotheses are then sent to the pose refiner, and from here, the implementation follows the original FoundationPose.
Scaling object meshes. The meshes output by InstantMesh are of arbitrary scale. To deal with this, we use the estimated depth from monocular depth estimation models to backproject each pixel within the object mask, creating a point cloud of that object. To remove outliers in the point cloud, we use Open3D’s remove_statistical_outlier and cluster_dbscan methods. Next, we find the axis-aligned bounding box of the point cloud and mesh. We calculate a scale factor $s=\dfrac{p}{m}$ where $p$ and $m$ are the diagonal lengths of the point-cloud and mesh bounding boxes respectively. Finally, we uniformly apply $s$ to the generated mesh to get an approximation of the true mesh scale.
ORION RGB vs RGB-D implementation. Below we summarize the implementation-level differences between our RGB-D and RGB-only pipelines. These differences affected how object correspondence and end-effector (EE) trajectories are obtained, and which assumptions are introduced in the pipeline
  • Tracking and transform estimation. Our RGB-D pipeline tracks 3D keypoint trajectories on objects and uses global registration of object point clouds to estimate the correspondence transform between demonstration and rollout scenes:
    $T_{\hat{o}\rightarrow o}\;=\;\operatorname{GLOBAL\_REGISTRATION}(\hat{o},o),$
    where $\hat{o}$ and $o$ are object point clouds of the demonstration and rollout objects, respectively.
    The RGB-only pipeline directly estimates the full 6-DOF pose of the target object in each frame during the demonstration, along with the initial pose of the target and reference objects during rollout. Thus, we only need to find the desired end pose of the target object during rollout. In this case, o and $o\text${ and $\hat${o} refers to the object’s pose. We find the rollout target end pose with
    ${o}_{target}^{end}=o_{ref}^{start}\cdot(\hat{o}_{ref}^{start})^{-1}\cdot\hat{o}_{target}^{start}$
    removing the need for an intermediate point-cloud registration step
  • End-effector (EE) pose sequence. The RGB-D pipeline finds the EE sequence by solving the optimization problem in 1. However, in the RGB-only case, because we already have the pose sequence of the object, we can apply a single transform to all elements of $o_{target}$ to retrieve $\{T\}$
  • Grasp selection. The RGB-only pipeline chooses grasp locations using object geometry heuristics derived from estimated meshes, rather than attempting to recover demonstrated hand/grasps from video. This is an assumption that simplifies grasp selection but can bias results relative to the RGB-D pipeline, which attempts to infer demonstrated hand information. In addition, the pose of this grasp selection is the transform applied to all elements of $o_{target}$ to retrieve the EE pose trajectory.
Object‑Centric Trajectory Warping
Algorithm 3 Translation Warping
1:3D demonstration object trajectory $\tau$, Desired 3D new start and end points $s,e$.
2:Warped 3D object trajectory $\tau^{warped}$
3:// Normalize trajectory and calculate rotation with Rodrigues formula
[…]
6:// Calculate rotation with Rodrigues formula
[…]
9:// Transform back to workspace frame
[…]
11:
12:return $\tau^{warped}$
Algorithm 4 Rotation Warping
1:6-DOF demonstration object trajectory $\tau$, desired 6-DOF start $s$ and end $e$. For some 4x4 6-DOF pose matrix $X$, we refer to the rotation as $Rot(X)\leftarrow X_{1:3,1:3}$ and translation as $t(X)\leftarrow X_{1:3,4}$.
2:Warped object rotation trajectory $R^{warped}$
[…]
5:// Rotate $R$ to align with demonstration direction
[…]
9:// Calculate boundary rotations and apply spherical linear interpolation
[…]
12:// Combine interpolated rotations with original trajectory
[…]
14:return $R^{warped}$
@@ Appendix B System Setup
Details of camera observations. As mentioned in Section 4, we use an iPad with a TrueDepth camera for collecting RGB-D human video demonstrations. We use an iOS app, Record3D, that allows us to access the depth images from the TrueDepth camera. We record RGB and depth image frames in sizes $1920\times 1080$ and $640\times 480$, respectively. To align the RGB images with the depth data, we resize the RGB frames to the size $640\times 480$. The app also automatically records the camera intrinsics of the iPhone camera so that the back-projection of point clouds is made possible. In the RGB-only case, we use the same videos with the depth channel removed, videos generated from generative models, and YouTube videos.
@@ Appendix C Success conditions of tasks
  • Peas-on-plate: The peas can is placed upright in the large area of the plate.
  • Pour-juice: The juice box is tilted 45 degrees in the direction of its opening, such that the opening ends above a container.
  • Cheese-on-plate: The cream cheese box is placed in the large area of the plate.
@@ Appendix D Additional Details on Experiments
Appendix E Prompts sent to VLM
[…]
A description of the video from which the demonstration frames were taken is ”demo_video_description”
Your output format is:
"similar_obj_names": [NAME_OF_SIMILAR_OBJ_1,
                    NAME_OF_SIMILAR_OBJ_2, ...]
"explanation": "Here, you will explain your reasoning for
                choosing similar_obj_names[i] as the
                similar object to obj_names[i]"
Example:
"similar_obj_names": [red water bottle, checkered plate]
"explanation": "obj_names[0] is ’green thermos’, and the
closest related object in the image is a red water bottle,
because they are both drinking objects which are both standing
upright, are both cylindrical, and both around 0.3 meters.
obj_names[1] is ’purple plate’, and the closest related
object in the image is a checkered plate, which has the
same shape and function as the purple plate, and can be
used as a reference object for the red water bottle.
Taking the whole real-world scene into consideration,
we can place the red water bottle onto the checkered
plate, just as the green thermos is placed onto
the purple plate in the demonstration video."
[…]
Identify task relevant objects, then classify into target and reference objects. The prompts used here are largely similar to the ones used in OKAMI [51].
[…]
Find task relevant objects prompt: You need to analyze what the human is doing in the images, then tell me: 1. All the objects in front of the scene (mostly on a plane similar to a table). You should ignore the background objects. 2. The objects of interest. They should be a subset of your answer to the first question. They are likely the objects manipulated by the human or objects that interact with other objects. Objects that the human uses (like utensils) to directly manipulate other objects should NOT be included as task relevant objects. Note that there are irrelevant objects in the scene, such as objects that are not interacted with by the human and/or the other objects. You should ignore the irelevant objects. Make sure to include a color and texture description of the relevant objects. This is especially important if there are many instances of a single type of object. For example, if there are blue, green, and yellow balls, be sure to specify each one individually. Do NOT just say ”colorful balls”. Finally, if any objects emit liquid, do not include the liquid itself as an object you return. Your output format is:
    "video_description": "The human is xxx.",
    "all_objects": ["OBJECT1", "OBJECT2", ...],
    "objects_of_interest": ["OBJECT1", "OBJECT3", ...],
Ensure the response can be parsed by Python ’json.loads’, e.g.: no trailing commas, no single quotes, etc.
Example:
    "video_description": "The human is picking up
    a red ball with blue dots and putting it in
    a green cup",
    "all_objects": ["blue ball with stripes",
                    "small green ball",
                    "red ball with blue dots",
                    "green cup"],
    "objects_of_interest": ["red ball with blue dots",
                            "green cup"],
[…]
Identify target objects prompt: The following images shows a manipulation motion, where the human is manipulating an object or objects.
Your task is to determine which objects are being manipulated in the images below. You need to choose from the following objects: {task_relevant_object_candidates}.
Tips: A manipulated object is an object that the human is interacting with, such as picking up, moving, or pressing, and it is in contact with the human or an object (like a utensil) that the human is using. If there are multiple objects of the same type, be sure to specify the color, texture, etc. that correspond to the manipulated object
Your output format is:
"manipulate_object_names": [
            "MANIPULATE_OBJECT_DESCRIPTIVE_NAME_1",
            "MANIPULATE_OBJECT_DESCRIPTIVE_NAME_2",
Example:
    "manipulate_object_names": ["red ball with blue dots",
                                "green cup"],
[…]
Identify reference objects prompt: The following images shows a manipulation motion, where the human is manipulating the objects in the list - {manipulate_obj_names}.
Please identify the reference object in the image below, which could be an object on which to place an object in the list {manipulate_obj_names}, or an object that something in the list {manipulate_obj_names} is interacting with. You need to first identify whether there is a reference object. If so, you need to output the reference object’s name chosen from the following objects: {task_relevant_object_candidates}. Note that there may not necessarily be a reference object. Examples of reference objects are: a coaster, if the motion involves picking up a cup and placing it on a coaster; a bowl, if the motion involves picking up a bottle and pouring it into the bowl.
Finally, make sure your proposed ith reference object corresponds to the given ith manipulated object.
"reference_object_names": ["REFERENCE_OBJECT_NAME_1",
                          "REFERENCE_OBJECT_NAME_2", ...]
                            or ["None"],
"reasoning": "Here, you will explain your reasoning for
choosing these objects as the reference object,
or for saying that there are no reference objects"
Your output format is:
Example:
    "reference_object_names": ["white circular plate",
                                "brown coaster"],
    "reasoning": The white circular plate is a reference
    object because the human is placing the red ball on it.
    The brown coaster is a reference object because
    the human is placing the green cup on it.
[…]